OpenStudy (areca):

3x + y = 1 y + 4 = 5x

OpenStudy (mathmale):

Please always post the instructions for problems.

OpenStudy (jiteshmeghwal9):

3x+y=1 y=1-3x...(1) substitute eqn (1) in the equation y+4=5x to gt x